# Imaging of I Zw 18 by JWST: I. Strategy and First Results of Dusty Stellar Populations

### v1

@@ 5 Summary
We present for the first time JWST near- and mid-IR imaging data using NIRCam and MIRI of the XMP SF dwarf galaxy I Zw 18. The extremely low levels of heavy-element enrichment and high level of active star formation make it an excellent accessible analog for the types of small systems which are expected to have been ubiquitous in the very early universe, and acted as major contributors to its overall chemical enrichment and reionization. With eight wide-band filters combining NIRCam (F115W, F200W, F356W, and F444W) and MIRI (F770W, F1000W, F1500W, F1800W), we show deep, high angular resolution images at IR wavelengths comparable to that of optical-band HST data, and improved by over an order of magnitude over that of existing mid-IR Spitzer images. Our NIRCam data clearly show extreme populations of bright, recently-formed massive stars, located predominantly among two central lobes of star formation (the NW and SE components). In addition, our MIRI imaging data demonstrates the bulk mid-IR emission characteristics of I Zw 18 at previously-unattainable resolution, exhibiting predominantly wispy and clumpy nature in the NW and SE components, respectively.
@@
Utilizing the four NIRCam filters, we have constructed CMDs and performed color-cut analyses from which we identified candidate populations of dusty evolved stars (RSGs and AGB stars) in this galaxy, alongside bright YSOs. We employ a CMD utilizing the shortest-wavelength NIRCam filters (F200W vs. F115W – F200W), which demonstrate the greatest point-source sensitivity and therefore possesses the largest number of sources, to analyze these populations. Having estimated the location of the TRGB, we divide those sources which lie above it (F200W 25.6) into left- and right-branch populations, pertaining to younger (UMS) and older (evolved) stars, respectively. Harnessing next a CMD with wider color baseline (F444W vs. F115W – F444W), we identify subsets of dust-enshrouded source candidates as those found above the TRGB and with color values 0.5, many of which are expected to be C-rich TP-AGB stars, but which also include O-rich AGB stars, RSGs, and bright YSOs.
@@
Further investigation of YSO or YSO cluster candidates was accomplished via the F200W – F444W vs. F115W vs. F200W CCD, building upon techniques implemented for recent JWST study of the NGC 346 SF region in the SMC. A small population of unresolved sources which satisfy the criteria F115W – F200W 1.2 and F200W – F444W 1.0 were isolated, with their locations in the galaxy correlated with MIRI F770W emission in the SE and NW lobes of star formation. Our findings suggest that the prevailing demographics of I Zw 18’s NW and SE star-formation regions reflect that of younger and older stellar populations, respectively, consistent with a staggered SFH.
@@
Point sources affiliated with I Zw 18’s companion system, Component C, were also extracted. A similar analysis was performed, finding CMD structures closely matching those constructed from the galaxy’s main body. We find a comparative lack of detections in the CMD location representative of Component C’s young, hot star population, suggesting that the bulk of its most recent star-formation episode took place longer ago than that of I Zw 18’s main body, consistent with the literature. In addition, the small number of detected dust-enshrouded source candidates are preferentially located in the bright end of the evolved star branch.

### v2

@@ 5 Summary
We present for the first time JWST near- and mid-IR imaging data using NIRCam and MIRI of the XMP SF dwarf galaxy I Zw 18. The extremely low levels of heavy-element enrichment and high level of active star formation make it an excellent accessible analog for the types of small systems which are expected to have been ubiquitous in the very early Universe, and acted as major contributors to its overall chemical enrichment and reionization. With eight wide-band filters combining NIRCam (F115W, F200W, F356W, and F444W) and MIRI (F770W, F1000W, F1500W, and F1800W), we show deep, high-angular-resolution images at IR wavelengths comparable to that of optical-band HST data, and improved by over an order of magnitude over that of existing mid-IR Spitzer images. Our NIRCam data clearly show extreme populations of bright, recently-formed massive stars, located predominantly among two central lobes of star formation (the NW and SE components). In addition, our MIRI imaging data demonstrate the bulk mid-IR emission characteristics of I Zw 18 at previously unattainable resolution, exhibiting a predominantly wispy and clumpy nature in the NW and SE components, respectively.
@@
Utilizing the four NIRCam filters, we have constructed CMDs and performed color-cut analyses from which we identified candidate populations of dusty evolved stars (RSGs and AGB stars) in this galaxy, alongside bright YSOs. We employ a CMD utilizing the shortest-wavelength NIRCam filters (F200W vs. F115W–F200W), which demonstrate the greatest point-source sensitivity and therefore possess the largest number of sources, to analyze these populations. Having estimated the location of the TRGB, we divide those sources which lie above it (F200W 25.6) into left- and right-branch populations, pertaining to younger (UMS) and older (evolved) stars, respectively. Harnessing next a CMD with wider color baseline (F444W vs. F115W–F444W), we identify subsets of dust-enshrouded source candidates as those found above the TRGB and with color values 0.5, many of which are expected to be C-rich TP-AGB stars, but which also include O-rich AGB stars, RSGs, and bright YSOs.
@@
Further investigation of YSO or YSO cluster candidates was accomplished via the F200W–F444W vs. F115W–F200W CCD, building upon techniques implemented for recent JWST study of the NGC 346 SF region in the SMC. A small population of unresolved sources which satisfy the criteria F115W–F200W 1.2 and F200W–F444W 1.0 were isolated, with their locations in the galaxy correlated with MIRI F770W emission in the SE and NW lobes of star formation. Our findings suggest that the prevailing demographics of I Zw 18’s NW and SE star-formation regions reflect that of younger and older stellar populations, respectively, consistent with a staggered SFH.
@@
Point sources affiliated with I Zw 18’s companion system, Component C, were also extracted. A similar analysis was performed, finding CMD structures closely matching those constructed from the galaxy’s main body. We find a comparative lack of detections in the CMD location representative of Component C’s young, hot star population, suggesting that the bulk of its most recent star-formation episode took place longer ago than that of I Zw 18’s main body, consistent with the literature. In addition, the small number of detected dust-enshrouded source candidates are preferentially located in the bright end of the evolved stars branch.